OpenStudy (anonymous):

What is an equation of the circle whose center is ( -3,1) and whose radius is 10 ? a. (x+3)^2 + ( y-1)^2 = 10 b. (x+3) ^2 + ( y-1)^2 = 100 c. ( x-3)^2 + (y+1)^2 = 10 d. (x-3)^2 + (y+1)^2 = 100

OpenStudy (anonymous):

center (h,k) (x-h)+(y-k)=r^2

OpenStudy (anonymous):

i dont understand .... whats h and k ? @LagrangeSon678

OpenStudy (anonymous):

h and k is the center. In your case, your h is the x coordinate, -3. your k is the y coordinate, 1

OpenStudy (anonymous):

just replace h with -3 and k with 1 in the equation of the circle:(x-h)^2+(y-k)^2=r^2